# and quantum information science

### v1

@@ 6.3.3 SYK model
The authors of [16] originally suggested Krylov complexity as a diagnostic of chaos. They proposed that this quantity would grow exponentially in chaotic systems because of the rapid spreading of operators expected there. In particular, according to [16], if the Lanczos coeffcients grow linearly at large $n$
@@
the Krylov complexity will grow exponentially at early times
@@
The first inequality was proven at infinite temperature in [16] and under the assumption of certain analytic and smoothness properties of the Lanczos coefficients, the right inequality was proven at finite temperature in [157, 158]. The inequality $\lambda_{\text{OTOC}}\leq\frac{2\pi}{\beta}$ where $\beta$ is the inverse temperature $\beta=1/T$ is the bound on chaos in quantum mechanical systems from [5] (the definition of $\lambda_{\text{OTOC}}$ is reviewed in Sec. 2).
@@
The SYK model admits some simplifications in the limit of infinite $N$, and even more when we take $q$ to also be large.202020 Here we mean that we first take the limit $N\rightarrow\infty$ and then we take $q\rightarrow\infty$. In the language of double scaled SYK (DSSYK), where $N$ is taken to infinity keeping the ratio $\lambda_{\text{SYK}}=2q^{2}/N$ fixed, our limit has $\lambda_{\text{SYK}}=0$. In this case, the survival amplitude averaged over the SYK ensemble is known analytically and one can use it to extract an averaged Lanczos sequence. At large $N$, some quantities are self averaging including the thermal two-point function [164] and many texts including the pioneering [16] work under the assumption that the Krylov complexity is similarly self-averaging. These works extract the Lanczos sequence from the ensemble averaged survival amplitude using the moment method described in Sec. 6.1.2. As we will see in below in the discussion of Random Matrix Theory in Sec. 6.3.4, ensemble averages of this kind match the coarse-grained Lanczos coefficients described in Sec. 6.1.4, but do not capture higher moments such as covariances of the Lanczos spectrum. Applying the moment method to the ensemble averaged survival amplitude the leading contribution to the Lanczos coefficients at large $q$ reads
@@
In this limit, the leading large $q$ contribution to the Krylov exponent is exactly equal to the Lyapunov exponent [16]: [^21] A particularly interesting case is the limit of low temperatures (large $\beta$). In this case both exponents approach the chaos bound $2\pi/\beta$.
@@
This surprising equality suggested that the Krylov exponent, which is simpler to compute that the OTOC, would provide a shortcut for computing the Lyapunov exponent of the latter. However, at large but finite $q$, the equality fails at order $1/q$, that is [165]
@@
The perturbation is an infrared deformation of the original SYK since it has effective dimension $\tilde{q}/q<1$ around the fixed point of the original Hamiltonian. This means that as we keep flowing to the infrared a new fixed point will appear that is dominated by the deformation. This type of deformation has been studied in, e.g., [166, 167, 168, 169, 170, 171, 172, 173, 174, 175]. In the context of chaos, we expect the infrared deformation to induce a transition between two chaotic regimes, or if $\tilde{q}=2$, between chaos and integrability along the RG flow controlled by the temperature parameter. We can try to diagnose these chaos-chaos or chaos-integrability transitions using the OTOC and Krylov exponents [165]. Fig. 8 show that while the Lyapunov exponent exhibits features indicative of such transitions, the Krylov exponent does not.
@@
Perhaps this should not come as a surprise, as the starting point for extracting the Lanczos sequence in the above discussion was the ensemble averaged two point function and Lanczos coefficients. As the next section suggests, to diagnose chaos we may need the ensemble fluctuations and correlations of the Lanczos sequence rather than their average. Additionally, the studies described above are investigating early time chaos. As we will discuss below, the spread complexity of chaotic systems grows to a peak at late times that are exponential in the entropy, and then slopes back down to a plateau. The presence of this peak and slope depend on characteristic spectral correlations that are absent in integrable theories with Poisson spectra [17]. We will discuss this below in the context of Random Matrix Theory, but all the characteristic behavior of spread complexity that we will describe applies equally to chaotic SYK models.
@@ 6.3.4 Random Matrix Theory
$\frac{1}{Z_{\beta,N}}e^{-\frac{\beta N}{4}\mathrm{Tr}[V(H)]}\,,$ (91)
where $H$ is the Hamiltonian, $V(H)=\sum_{n=0}^{\infty}v_{n}\mathrm{Tr}(H^{n})$ is a potential, and the partition sum $Z_{\beta,N}$ normalizes the measure. The Dyson index is $\beta=1,2,4$ depending on whether we choose $H$ to be Hermitian, real and symmetric, or quaternionic, corresponding to unitary, orthogonal, or symplectic ensembles. If $V$ is quadratic, we say that the Random Matrix Theory (RMT)is described by the Gaussian Unitary, Orthogonal or Symplectic Ensembles (GUE,GOE, GSE). We are going to study time evolution and spread complexity of a generic state in the basis in which the Hamiltonian is drawn. Since the matrix is being drawn randomly we can pick this to be the state $(1,0,0,\cdots)$ if we like.
Suppose we diagonalize the Hamiltonian in the exponent of (91) by a unitary transformation. Famously, the change of variables introduces the Vandermode determinant $\Delta=\prod_{i}<j|\lambda_{i}-\lambda_{j}|^{\beta}$ in the measure, leading to a distribution over eigenvalues of the form
$p(\lambda_{1},\cdots\lambda_{N})=Z_{\beta,N}e^{-\frac{\beta N}{4}\mathrm{Tr}[V(\Lambda)]}\prod_{i<j}|\lambda_{i}-\lambda_{j}|^{\beta}=Z_{\beta,N}e^{-\frac{\beta N}{4}\sum_{n}v_{n}\sum_{k}\lambda_{k}^{n}}\prod_{i<j}|\lambda_{i}-\lambda_{j}|^{\beta}\,,$ (92)
where $V(\Lambda)$ is the potential evaluated on the diagonal matrix of eigenvalues. There is also an important relation between the potential $V$ and the ensemble averaged density of states:
@@
$p(a_{0},\cdots a_{N-1};b_{0},\cdots b_{N-1})\propto\left(\prod_{n=1}^{N-1}b_{n}^{(N-n)\beta-1}\right)e^{-\frac{\beta N}{4}Tr[V(H)]}\,,$ (94)
where within the trace $H$ is understood as tridiagonalized to the Lanczos form.
At large $N$ the distribution (94) is sharply peaked. So we can use the method of saddle points to extract the ensemble average (the saddle point) and the covariances (from the quadratic expansion around the saddle point). The ensemble average precisely reproduces (61) which we already derived through other means, and which applies to any theory for which we know the density of states. A formula for the covariances of the $a_{n}$ and $b_{n}$ is derived in [126, 156]. We will not repeat the derivation here for lack of space, but the result is
@@
Following [17] the spread complexity of the time-evolved thermofield-double state for a GUE Hamiltonian is plotted in Fig. 9 for various matrix sizes $N$ and inverse temperatures $\beta$. The dynamics exhibit four characteristic regimes: a linear ramp up to a peak that is exponential in the entropy, followed by a slope down to a plateau. The authors of [17] showed that the peak and slope arise from eigenvalue correlations known as spectral rigidity in chaotic theories. These correlations can be removed by drawing eigenvalues at random from the Wigner semicircle distribution and the evolution with such a spectrum (Fig. 9 in light hues) shows only the ramp and plateau. Thus, the peak and slope are characteristic of late time chaos. The same patterns appear in the late time spread complexity of the chaotic SYK model [17].
Figure 9: Quantum state complexity of the time-evolved thermofield-double state over an exponentially large period of time for different values of $N$ and $\beta$, as described in the main text. Dark Hues: GUE ensemble. Going from highest (blue) to lowest (yellow) curves we have $\beta=\{0,1,2,5,10\}$. In each case we have plotted ensembles with $N=\{1024,1280,1536,1792,2048,2560,3072,3584,4096\}$. Complexity grows linearly to a peak, followed by a downward slope to a plateau. Light Hues: Ensemble with the same density of states as GUE, but without correlations between eigenvalues. In this case, the curves plateau without reaching a peak followed by a downward slope.
@@
where $X_{i}$, $Y_{i}$, $Z_{i}$ are Pauli spin operators on site $i$; $A_{k}$, $B_{k}$ and $C_{k}$ are numerical parameters, and $H_{\text{site}}=0.5Z_{3}+0.3X_{3}$ is a one-site disorder operator that helps breaks parity and symmetry under the action of $\sum_{i}Z_{i}$. The model defined by $H_{\text{int1}}$ is integrable, while $H_{\text{cha1}}$ is chaotic as suggested by their level statistics, and have been explored in the literature [178, 155]. Indeed, as shown in [156], the analytical formulae developed above give an excellent description of the Lanczos coefficients in all cases, but the variances and covariances deviate significantly from the RMT formula for integrable spin chains, and in particular are much larger.
We can also write down an explicit formula for the height of the late time plateau in the spread complexity. First, let $\hat{n}=\sum_{n}n\lvert K_{n}\rangle\langle K_{n}\rvert$ be the position operator along the Krylov chain associated to an initial state $\lvert\phi\rangle$. Then, define the complexity of an energy eigenstate $\lvert E\rangle$ relative to $\lvert\phi\rangle$ as $\mathcal{C}_{E}=\langle E\rvert\hat{n}\lvert E\rangle$. A short calculation [156] shows that the height of the late time plateau in the spread complexity will be
@@
where $p(E)=|\langle E|\phi\rangle|^{2}$ is the initial probability of eigenstate $\lvert E\rangle$. If $p(E)$ is continuous we can replace $\mathcal{C}_{E}$ by the local mean to get
@@ 6.4 Open problems and directions
We will conclude this section on the complexity of the spread of states and operators with a discussion of some open questions and progress towards answering them.
@@
These quantities are known exactly in some QFTs. For example, in 2D CFTs, thermal two-point correlators are universally determined by Ward identities and are the same for both integrable and chaotic CFTs. In this case, the there is an infinite sequence of $b_{n}$ Lanczos coefficients that grow linearly for large $n$ [185]. At least at first sight, this serves as a counterexample to the universal operator growth hypothesis. This issue was explored in [186, 187], where the authors argued that to meaningfully characterize operator growth in QFTs, one must first introduce appropriate cut-offs. They analyzed the effects of both UV and IR cut-offs, and concluded that the chaotic or scrambling properties of QFTs may be captured by $b_{n}$’s larger than the UV cut-off $\Lambda$. Meanwhile, the IR cut-off (e.g. modeled by introducing a mass) induced a separation (staggering) between even and odd Lanczos coefficients that grew linearly with different intercepts [187].
Thus, distinguishing characteristic features of chaotic and integrable operator growth in continuum QFTs remains an open problem. A promising approach could be to employ von Neumann algebras and algebraic QFTs. Some progress in this direction has been made in [188, 189], but a direct connection with Krylov-basis methods has not been established. Alternatively, we could compute the covariances of the Lanczos coefficients in QFT, in view of the conjecture in [126, 156] that their structure separates chaos from integrability. In theories with a discrete spectrum, it may also help to project onto a microcanonical band of energies (see also [190]) thereby rendering the accessible Hilbert space finite dimensional, thus enabling us to directly adopt the ideas and methods that have been developed for such systems.
@@
The relation between Krylov and Nielsen’s approaches has been debated in several works. In [191], the authors argued that Krylov complexity can be interpreted geometrically as a distance in a certain projection of the Fubini-Study geometry that is natural in quantum optics, making it amenable to a circuit interpretation. However, as we already saw from the symmetry based examples, Krylov complexity, as a function of time, is proportional to the volume in the Fubini-Study geometry. As a volume, rather than a length, it does not respect some properties of geodesics, such as the triangle inequality. Such properties were suggested in [192] as an obstruction to Krylov complexity serving as a distance measure. However, it is not clear that the Fubini-Study setting is the one in which we should seek a relation with Nielsen complexity. Indeed, [193] argued that, in certain setups, there can be a relation between the time average of spread complexity of state evolution and an upper bound on Nielsen complexity with specific fine-tuned penalty factors. Recalling that the Nielsen approach has ambiguities in the choice of gate set and cost function, it remains possible that some judicious choice could establish a direct relation between the Nielsen and Krylov/spread approaches to complexity, at least for some specific choices of states. Perhaps useful progress could be achieved by formulating Krylov/spread complexity approach to general, time-dependent Hamiltonians (see progress in [194]). This would be closer in spirit to the starting point in Nielsen’s definitions and could allow us to use bounds on circuit complexity to constrain Krylov complexity. Indeed, if we treat the evolution parameter as a circuit time, we can mathematically think about the complexity of formation of states in the language of spread complexity [195]. Thus it remains an interesting problem to relate the complexity of state/operator spreading to circuit complexity.
@@
It would be interesting to extend the spread/Krylov complexity approaches to new systems. First, the generalization to time dependent Hamiltonians is important to achieve for the reasons described above, and because we would like to understand systems that are driven by external sources. Second, if we contemplate external sources, we should also consider open quantum systems, whose dynamics are controlled by Lindbladian evolution of the density matrix. There is a substantial literature on the Linbladian formalism which is partly reviewed in [3]. The interaction with an external bath with an open system certainly affects the structure of entanglement spreading (see, e.g., [196]), and has generally been a subject of significant recent interest in condensed matter physics through the study of measurement driven phase transitions (see, e.g., [197]). Another interesting avenue is to study the spread complexity of Carrollian theories of relevance for flat space holography. Likewise it would be interesting to study the kind of chiral theories that appear in the AdS/CFT dual to highly rotating extremal black holes [198, 199, 200] and the matrix models that appear as the dual to the 11-dimensional M-theory [201, 202, 203].

### v2

@@ 6.3.3 SYK model
The authors of [21], originally suggested Krylov complexity as a diagnostic of chaos. They proposed that this quantity would grow exponentially in chaotic systems because of the rapid spreading of operators expected there. In particular, according to [21], if the Lanczos coeffcients grow linearly at large $n$
@@
the Krylov complexity will grow exponentially at early times [^23] By early, we mean up to times of the order of the logarithm of the number of the degrees of freedom in the system (or forever, if the system has infinitely many degrees of freedom).
@@
The first inequality was proven at infinite temperature in [21] and, under the assumption of certain analytic and smoothness properties of the Lanczos coefficients, the right inequality was proven at finite temperature in [176, 177]. The inequality $\lambda_{\text{OTOC}}\leq\frac{2\pi}{\beta}$, where $\beta=1/T$ is the inverse temperature, is the bound on chaos in quantum mechanical systems from [10] (the definition of $\lambda_{\text{OTOC}}$ is reviewed in Sec. 2).
@@
The SYK model admits some simplifications in the limit of infinite $N$, and even more when we take $q$ to also be large.242424 Here we mean that we first take the limit $N\rightarrow\infty$ and then we take $q\rightarrow\infty$. In the language of DSSYK, where $N$ is taken to infinity keeping the ratio $\lambda_{\text{SYK}}=2q^{2}/N$ fixed, our limit has $\lambda_{\text{SYK}}\rightarrow 0$. In this case, the survival amplitude averaged over the SYK ensemble is known analytically and one can use it to extract an averaged Lanczos sequence. At large $N$, some quantities are self averaging including the thermal two-point function [183] and many texts, including the pioneering Ref. [21], work under the assumption that the Krylov complexity is similarly self-averaging. These works extract the Lanczos sequence from the ensemble averaged survival amplitude using the moment method described in Sec. 6.1.2.252525Sometimes the survival amplitude is referred to as the auto-correlation function or thermal two-point function in this context. As we will see below in the discussion of Random Matrix Theory in Sec. 6.3.4, ensemble averages of this kind match the coarse-grained Lanczos coefficients described in Sec. 6.1.4, but do not capture higher moments such as covariances of the Lanczos spectrum. Applying the moment method to the ensemble averaged survival amplitude, the leading contribution to the Lanczos coefficients at large $q$ reads
@@
In this limit, the leading large $q$ contribution to the Krylov exponent is exactly equal to the Lyapunov exponent for the OTOC correlators [21]:
@@
A particularly interesting case is the limit of low temperatures (large $\beta$). In this case both exponents approach the chaos bound $2\pi/\beta$. The surprising equality between the Krylov and Lyaponov exponent for the OTOC suggested that the Krylov exponent, which is simpler to compute than the OTOC, would provide a shortcut for computing the Lyapunov exponent of the latter. However, at large but finite $q$, the equality fails at order $1/q$, that is [184]
@@
The perturbation is an infrared deformation of the original SYK since it has effective dimension $\tilde{q}/q<1$ around the fixed point of the original Hamiltonian. This means that as we keep flowing to the infrared, a new fixed point will appear that is dominated by the deformation. This type of deformation has been studied in, e.g., [185, 186, 187, 188, 189, 190, 191, 192, 193, 194]. In the context of chaos, we expect the infrared deformation to induce a transition between two chaotic regimes, or if $\tilde{q}=2$, between chaos and integrability along the RG flow controlled by the temperature parameter. We can try to diagnose these chaos-chaos or chaos-integrability transitions using the OTOC and Krylov exponents [184]. Fig. 8 show that while the Lyapunov exponent exhibits features indicative of such transitions, the Krylov exponent does not.
@@
Perhaps this should not come as a surprise, as the starting point for extracting the Lanczos sequence in the above discussion was the ensemble averaged two point function and Lanczos coefficients. As the next section suggests, to diagnose chaos we may need the ensemble fluctuations and correlations of the Lanczos sequence rather than their average.
Additionally, the studies described above are investigating early time chaos. As we will discuss below, the spread complexity of chaotic systems grows to a peak at late times that are exponential in the entropy, and then slopes back down to a plateau. Equivalently, the peak occurs at a time linear in the dimension of the Hilbert space, which scales as the exponential of the entropy. The presence of this peak and slope depend on characteristic spectral correlations that are absent in integrable theories with Poisson spectra [22]. We will discuss this below in the context of Random Matrix Theory, but all the characteristic behavior of spread complexity that we will describe applies equally to chaotic SYK models.
@@ 6.3.4 Random Matrix Theory
$\frac{1}{Z_{\beta_{D},N}}e^{-\frac{\beta_{D}N}{4}\mathrm{Tr}[V(H)]}\,,$ (102)
where $H$ is a Hamiltonian matrix acting on an $N$–dimensional Hilbert space, $V(H)=\sum_{n=0}^{\infty}v_{n}\mathrm{Tr}(H^{n})$ is a potential, and the partition sum $Z_{\beta_{D},N}$ normalizes the measure. [^26] In this section we return to using $N$ for the Hilbert space dimension rather than the number of degrees of freedom. The Dyson index is $\beta_{D}=1,2,4$ depending on whether we choose $H$ to be Hermitian, real and symmetric, or quaternionic, corresponding to unitary, orthogonal, or symplectic ensembles. If $V$ is quadratic, we say that the Random Matrix Theory (RMT) is described by the Gaussian Unitary, Orthogonal or Symplectic Ensembles (GUE,GOE, GSE). We are going to study time evolution and spread complexity of a generic state in the basis in which the Hamiltonian is drawn. Since the matrix is being drawn randomly, we can pick this to be the state $(1,0,0,\cdots)$ without loss of generality.
Suppose we diagonalize the Hamiltonian in the exponent of (102) by a unitary transformation. Famously, the change of variables introduces the Vandermode determinant $\Delta=\prod_{i<j}|\lambda_{i}-\lambda_{j}|^{\beta_{D}}$ in the measure, leading to a distribution over eigenvalues of the form
$p(\lambda_{1},\cdots\lambda_{N})=Z_{\beta_{D},N}e^{-\frac{\beta_{D}N}{4}\mathrm{Tr}[V(\Lambda)]}\prod_{i<j}|\lambda_{i}-\lambda_{j}|^{\beta_{D}}=Z_{\beta_{D},N}e^{-\frac{\beta_{D}N}{4}\sum_{n}v_{n}\sum_{k}\lambda_{k}^{n}}\prod_{i<j}|\lambda_{i}-\lambda_{j}|^{\beta_{D}}\,,$ (103)
where $V(\Lambda)$ is the potential evaluated on the diagonal matrix of eigenvalues $\lambda_{k}$. There is also an important relation between the potential $V$ and the ensemble averaged density of states:
@@
$p(a_{0},\cdots a_{N-1};b_{0},\cdots b_{N-1})\propto\left(\prod_{n=1}^{N-1}b_{n}^{(N-n)\beta_{D}-1}\right)e^{-\frac{\beta_{D}N}{4}Tr[V(H)]}\,,$ (105)
where within the trace, $H$ is understood as tridiagonalized to the Lanczos form.
At large $N$ the distribution (105) is sharply peaked. So we can use the method of saddle points to extract the ensemble average (the saddle point) and the covariances (from the quadratic expansion around the saddle point). The ensemble average precisely reproduces (71), which we already derived through other means, and which applies to any theory for which we know the density of states. A formula for the covariances of the $a_{n}$ and $b_{n}$ is derived in [23, 24]. We will not repeat the derivation here for lack of space, but the result is
@@
Following [22], the spread complexity of the time-evolved thermofield-double state for a GUE Hamiltonian is plotted in Fig. 9 for various matrix sizes $N$ and inverse temperatures $\beta$. The dynamics exhibit four characteristic regimes: a linear ramp up to a peak that is exponential in the entropy (and hence linear in $N$, the Hilbert space dimension), followed by a slope down to a plateau. The authors of [22] showed that the peak and slope arise from eigenvalue correlations known as spectral rigidity in chaotic theories. These correlations can be removed by drawing eigenvalues at random from the Wigner semicircle distribution – this produces a spectrum without eigenvalue repulsion, like the Poisson spectra expected for integrable theories. The evolution with such a spectrum (Fig. 9 in light hues) shows only the ramp and plateau. Thus, the peak and slope are characteristic of late time chaos. The same patterns appear in the late time spread complexity of the chaotic SYK model [22].
Figure 9: Quantum state complexity of the time-evolved thermofield-double state over an exponentially large period of time for different values of $N$ and $\beta$, as described in the main text. Dark Hues: GUE ensemble. Going from highest (blue) to lowest (yellow) curves we have $\beta=\{0,1,2,5,10\}$. In each case, we have plotted ensembles with $N=\{1024,1280,1536,1792,2048,2560,3072,3584,4096\}$. Complexity grows linearly to a peak, followed by a downward slope to a plateau. Light Hues: Ensemble with the same density of states as GUE, but without correlations between eigenvalues. In this case, the curves plateau without reaching a peak followed by a downward slope. Figure taken from [22].
@@
where $X_{i}$, $Y_{i}$, $Z_{i}$ are Pauli spin operators on site $i$; $K$ is the number of spins; $A_{k}$, $B_{k}$ and $C_{k}$ are numerical parameters; and $H_{\text{site}}=0.5Z_{3}+0.3X_{3}$ is a one-site disorder operator that helps breaking parity and symmetry under the action of $\sum_{i}Z_{i}$. The model defined by $H_{\text{int1}}$ is integrable, while $H_{\text{cha1}}$ is chaotic as suggested by their level statistics, and have been explored in the literature [197, 175]. Indeed, as shown in [24], the analytical formulae developed above give an excellent description of the Lanczos coefficients in all cases, but the variances and covariances deviate significantly from the RMT formula for integrable spin chains, and in particular are much larger.
We can also write down an explicit formula for the height of the late time plateau in the spread complexity. First, let $\hat{n}=\sum_{n}n\lvert K_{n}\rangle\langle K_{n}\rvert$ be the position operator along the Krylov chain associated to an initial state $\lvert\phi\rangle$. Then, define the complexity of an energy eigenstate $\lvert E\rangle$ relative to $\lvert\phi\rangle$ as $\mathcal{C}_{E}=\langle E\rvert\hat{n}\lvert E\rangle$. A short calculation in [24] shows that the height of the late time plateau in the spread complexity will be
@@
where $p(E)=|\langle E|\phi\rangle|^{2}$ is the initial probability of eigenstate $\lvert E\rangle$. If $p(E)$ is continuous, we can replace $\mathcal{C}_{E}$ by the local average in a small window over energies to get a coarse-grained eigenstate complexity $\overline{\mathcal{C}_{E}}$. A calculation in [24] shows that $\overline{\mathcal{C}_{E}}$ is entirely determined by Lanczos coefficients, which are in turn determined in terms of the density of states following Eq. (71). In terms of the coarse grained eigenstate complexities, and in the large system limit, we can take a continuum limit of (109), which yields
@@ 6.4 Open problems and directions
We conclude this section on the complexity of the spread of states and operators with a discussion of some open questions and progress towards answering them.
@@
These quantities are known exactly in some QFTs. For example, in 2D CFTs, thermal two-point correlators are universally determined by Ward identities and are the same for both integrable and chaotic CFTs. In this case, there is an infinite sequence of $b_{n}$ Lanczos coefficients that grow linearly for large $n$ [204]. At least at first sight, this serves as a counterexample to the universal operator growth hypothesis. This issue was explored in [205, 206], where the authors argued that to meaningfully characterize operator growth in QFTs, one must first introduce appropriate cut-offs. They analyzed the effects of both UV and IR cut-offs, and concluded that the chaotic or scrambling properties of QFTs may be captured by $b_{n}$’s larger than the UV cut-off $\Lambda$. Meanwhile, the IR cut-off (e.g. modeled by introducing a mass) induced a separation (staggering) between even and odd Lanczos coefficients that grew linearly with different intercepts [206].
Thus, distinguishing characteristic features of chaotic and integrable operator growth in continuum QFTs remains an open problem. A promising approach could be to employ von Neumann algebras and algebraic QFTs. Some progress in this direction has been made in [207, 208], but a direct connection with Krylov-basis methods has not been established. Alternatively, we could compute the covariances of the Lanczos coefficients in QFT, in view of the conjecture in [23, 24] that their structure separates chaos from integrability. In theories with a discrete spectrum, it may also help to project onto a microcanonical band of energies (see also [209]) thereby rendering the accessible Hilbert space finite dimensional, thus enabling us to directly adopt the ideas and methods that have been developed for the latter.
@@
The relation between Krylov and Nielsen’s approaches has been debated in several works. In [210], the authors argued that Krylov complexity can be interpreted geometrically as a distance in a certain projection of the Fubini-Study geometry that is natural in quantum optics, making it amenable to a circuit complexity interpretation. However, as we already saw from the symmetry-based examples, Krylov complexity, as a function of time, is proportional to the volume in the Fubini-Study geometry. As a volume, rather than a length, it does not respect some properties of geodesics, such as the triangle inequality. Such properties were suggested in [211] as an obstruction to Krylov complexity serving as a distance measure. However, it is not clear that the Fubini-Study setting is the one in which we should seek a relation with Nielsen’s complexity. Indeed, [212] argued that, in certain setups, there is a relation between the time average of spread complexity of state evolution and an upper bound on Nielsen’s complexity with specific fine-tuned penalty factors. Recalling that the Nielsen approach has ambiguities in the choice of basis of generators and cost function, it remains possible that some judicious choice could establish a direct relation between the Nielsen and Krylov/spread approaches to complexity, at least for some specific choices of states. Perhaps useful progress could be achieved by formulating Krylov/spread complexity approach to general, time-dependent Hamiltonians (see progress in [213]). This would be closer in spirit to the starting point in Nielsen’s definitions and could allow us to use bounds on circuit complexity to constrain Krylov complexity. Indeed, if we treat the evolution parameter as a circuit time, we can mathematically think about the complexity of formation of states in the language of spread complexity [214]. Thus it remains an interesting problem to systematically relate the complexity of state/operator spreading to circuit complexity.
@@
It would be interesting to extend the spread/Krylov complexity approaches to new systems. First, the generalization to time-dependent Hamiltonians is important to achieve for the reasons described above, and because we would like to understand systems that are driven by external sources. Second, if we contemplate external sources, we should also consider open quantum systems, whose dynamics are controlled by Lindbladian evolution of the density matrix [215]. There is a substantial literature on the Linbladian formalism which is partly reviewed in [3]. The interaction of an external bath with an open system certainly affects the structure of entanglement spreading (see, e.g., [216]), and has generally been a subject of significant recent interest in condensed matter physics through the study of measurement-driven phase transitions (see, e.g., [217]). The spread complexity of measurement-induced non-unitary dynamics has been explored in [160]. Meanwhile, [218] have uncovered various bounds and relations pertaining to the state/operator complexity of purifying a mixed state. Another interesting avenue is to study the spread complexity of Carrollian theories of relevance for flat space holography. Likewise it would be interesting to study the kind of chiral theories that appear in the AdS/CFT dual to highly rotating extremal black holes [219, 220, 221] and the matrix models that appear as the dual to the 11-dimensional M-theory [222, 223, 224]. In this direction, spread complexity has been studied in the planar limit of holographic theories in [225].
Applications to the physics of black holes
In Sec. 7.5 we will discuss how spread complexity of the double scaled SYK model is directly related to the length of the dual wormholes in two-dimensional Jackiw-Teitelboim gravity. But we could ask a broader question. Many lines of argument suggest that the microstates of black holes are immensely complex and may appear to be hidden behind a horizon simply because simple probes cannot discriminate between them – see, e.g., [226, 227, 228, 229]. Similarly, a number of authors have argued that in the AdS/CFT correspondence excitations behind the horizon are encoded in the CFT by exponentially complex operators [230, 231] in a non-isometric manner [232]. The information stored in Hawking radiation is also supposed to be encoded in an extremely complex manner, with the complexity expressed in the dual gravitational theory in terms of geometric obstructions preventing asymptotic observers from accessing the information [233]. At least in toy models of black hole evaporation like [234], such considerations can be explicitly realized and characterized in information-theoretic and complexity-theoretic terms [235]. But black holes can form from collapse of simple states like an ordered cloud of particles. So, the immense complexity of black hole microstates, and of the subsequent encoding of the initial state in the Hawking radiation, must be generated dynamically by evolution under the Hamiltonian of the theory. An important question is to understand how this happens. Perhaps the techniques of Krylov/spread complexity, or of Nielsen complexity from the previous section, can help. One recent result shows that if a holographic CFT time evolution operator is approximately pseudorandom or Haar random on a low energy subspace of states, there must be an event horizon in the dual theory of gravity [236]. Likewise, it seems that any region of AdS with sufficiently high holographic encoding complexity will be robustly protect from low-rank measurements, and hence be inaccessible to low-complexity observers [235]. It would be interesting to understand how the dynamical evolution of simple states can produce complex states with such properties.